OpenStudy (anonymous):

Help me please... Find the real or imaginary solutions of this solution of each equation by factoring. 8x^3-27=0 Using: a^3-b^3=(a-b)(a^2+ab+b^2)

OpenStudy (mertsj):

\[8x^3-27=(2x)^3-(3)^3\]

OpenStudy (mertsj):

In your formula, replace each a with 2x and each b with 3

OpenStudy (callisto):

8x^3-27=0 (2x)^3 - (3)^(3) = 0 Using the formula, (2x-3) ( 2x^2 + 6x +9) =0 (2x-3) = 0 or ( 2x^2 + 6x +9) =0 x=3/2 or 2[x^2 + 3x + (3/2)^2] - 2(3/2)^2 +9 =0 2(x + 3/2)^2 +27/2 = 0 2(x + 3/2)^2 = -27/2 (x + 3/2)^2 = -27/4 x+3/2 = (sqrt 27) i /2 or x+3/2 = - (sqrt 27) i /2 x= [(sqrt 27) -3] /2 or x = [-(sqrt 27)-3]/2

OpenStudy (anonymous):

how did you find the second part?

OpenStudy (callisto):

the second part?

OpenStudy (anonymous):

where is says x=3/2 from there on..

OpenStudy (callisto):

look at the (2x-3) = 0 first (2x-3) = 0 2x = 3 x=3/2, got it?

OpenStudy (anonymous):

ohhhhhh. yess

OpenStudy (callisto):

for ( 2x^2 + 6x +9) =0 part, you can either get the answer by completing square or using the quadratic formula

OpenStudy (callisto):

nah.. i guess i've made mistakes..

OpenStudy (callisto):

use the quad. formula, \[(-6\pm \sqrt{6^{2}-4(2)(9)})/[2(2)] = (-6\pm \sqrt{-36}) /4\]\[ =(-6 \pm 6i) /4 \] \[ =(-3 \pm 3i) /2 \]

OpenStudy (anonymous):

is there anything after that?